OpenStudy (anonymous):

Got a question from the "Tangent Planes and Differentials" section of multivariable calculus. Use differentials to estimate the amount of tin in a closed tin can with diameter 8cm and height 12cm if the tin is 0.04cm thick. The book gives and answer of 16cm^3 but I don't see how to do this one. Thanks!

OpenStudy (anonymous):

Volume of a cylinder(tin can) where r = radius = 4cm h = height = 12cm Applying the definition of the differential dz, The differential of V is: Suppose that the radius is decreased by 0.04 cm (dr = -0.04) and the height is decreased by 0,08 cm (dh = - 0.08). Then dV = 96 Pi (-0.04) + 16 Pi (-0.08) = 16.08 the amount of tin is approximately 16cm^3

OpenStudy (anonymous):

or this solution: dv = 2(pi)rhdr + (pi)r^2dh dr = .04 & dh = .08 dv = 16.08 cm^3

OpenStudy (anonymous):

Where did the 0.08 come from?

OpenStudy (anonymous):

from the given diameter dh=.08

OpenStudy (anonymous):

Oh wait, I see now. But dh should equal 0.04 as this is the thickness dimension. Also, since the can has both a top and a bottom the expression should be: \[dV=2 \pi rhdr + 2 \pi r^2dh=2 \pi(4)(12)(0.04) + 2 \pi (4)^2(0.04)\]This gives,\[dV=12.063+4.021\approx16cm^3\]Thanks for your help!

OpenStudy (anonymous):

yeah..thanks also and glad to help:-)